Consider the following analytical cell: Pt(s)|H2(g,0.10atm)|NaOH(aq,0.10M)||NaOH(aq,0.0010M)|Ag2O|Ag(s) a. If we assumed that all the reactants were under standard conditions, what is E°cell? b. What is Ecathode under the given conditions? c. What is Eanode under the given conditions? d. What is Ecell under the given conditions?
Answer #1

Pt(s)|H2(g,0.10atm)|NaOH(aq,0.10M)||NaOH(aq,0.0010M)|Ag2O|Ag(s)

the anodic reaction involves oxidation

H2(g) + 2OH-(aq) --->2H2O(l) + 2e-   E0anode = -0.83V

The cathodic reaction involves reduction

Ag2O(s) + H2O(l) + 2e --> 2Ag(s) + 2OH- E0cathode = +0.342V

E0cell = E0cathode - E0anode = 0.342 - (-0.83) = +1.172V

(b) Eanode = E0anode - 0.0592/2 log [OH-]^2 = -0.83 - 0.0296 (-6) = -0.6524V

(c) Ecathode = E0cathode - 0.0592/2 log (OH-)^2 = +0.342 +0.0592 = 0.4012 V

(d) Ecell = Ecathode - Eanode = 0.4012 - (-0.6524) = 1.0536 V